OpenStudy (anonymous):

evaluate the limit of 1/(x+1)^2 as x approaches -1

OpenStudy (anonymous):

limit of 1/(x+1)^2 as x approaches -1=\[limx->-1 \frac{ 1 }{ (x+1)^{2} }=\infty \]

OpenStudy (anonymous):



OpenStudy (anonymous):

ohhh okay thanks